For the following fuels burning in air, write down the stoichiometric reactions, the respective fuel–air ratios (F/A), and the heats of combustion per mole ( ¯qc) and per gm (qc). Assume the product water is in the gaseous state. Fuels: CO, H2, CH4, C7H16, C16H34, C6H6, CH3OH, C7H15OH. What can you say about the qcs for (a) CH4, C7H16, C16H34, and (b) C7H16, CH3OH, C7H15OH?
Answer #1

i) 2CO(g) + O2(g) \rightarrow 2CO2(g) (F/A = 1:2)

Heat of combustion (\DeltaHc) = -294 KJ/mol = -294/28 KJ/g = -10.5 KJ/g

ii) 2H2(g) + O2(g) \rightarrow 2H2O(g) (F/A = 2:1)

Heat of combustion (\DeltaHc) = -286 KJ/mol = -286/2 KJ/g = -143 KJ/g

iii) CH4(g) + 2O2(g) \rightarrow CO2(g) + 2H2O(g) (F/A = 1:2)

Heat of combustion (\DeltaHc) = -890 KJ/mol = -890/16 KJ/g = -55.625 KJ/g

iv) C7H16(l) + 11O2(g) \rightarrow 7CO2(g) + 8H2O(g) (F/A = 1:11)

Heat of combustion (\DeltaHc) = -4817 KJ/mol = -4817/100 KJ/g = -48.17 KJ/g

v) 2C16H34(l) + 49O2(g) \rightarrow 32CO2(g) + 34H2O(g) (F/A = 2:49)

Heat of combustion (\DeltaHc) = -10699 KJ/mol = -10699/226 KJ/g = -47.34 KJ/g

vi) C6H6(l) + 9O2(g) \rightarrow 6CO2(g) + 6H2O(g) (F/A = 1:9)

Heat of combustion (\DeltaHc) = -3271 KJ/mol = -3271/78 KJ/g = -41.94 KJ/g

vii) 2CH3OH(l) + 3O2(g) \rightarrow 2CO2(g) + 4H2O(g) (F/A = 2:3)

Heat of combustion (\DeltaHc) = -726 KJ/mol = -726/32 KJ/g = -22.69 KJ/g

viii) 2C7H15OH(l) + 21O2(g) \rightarrow 14CO2(g) + 16H2O(g) (F/A = 2:41)

Heat of combustion (\DeltaHc) = 4638 KJ/mol = 4638/116 KJ/g = 39.98 KJ/g

Note: Simple hydrocarbons release more heat than the alcohols in the corresponding combustion reaction. Higher alcohols absorb heat in their combustion.